10. Write an equation of a line in standard form for the given conditions. passes through (0, 0) and is perpendicular to the line 04x + 5y = 0
Answer #1

When 2 lines are perpendicular, the product of their slopes = -1. Therefore m1 x m2 = -1

The slope of line (5/4)x - y = 6 is m1 = 5/4

Therefore (5/4) x m2 = -1

m2 = -4/5.

This line passes through the point (0,0), therefore by the skope point method, we get

y - 0 = (-4/5) * (x - 0)

y = -(4/5)x

We need the equation of the line in the STandard Form = Ax + By + C = 0 form

Cross multiplying, 5y = -4x

Therefore 5y + 4x = 0 (Option 2)